A is 12 year old, B is 24 year ols. how many years ago was B's age 3 times that os A?

Answer by ewatrrr(24785)

x the number of years ago
24-x= 3(12-x)
2x = 12
x = 6 years ago
And..checking... 6 years ago ...
18 = 3*6